# svg_figures/

### v1

@@ Abstract
The increasing reliance on generative AI models has accelerated the generation rate of synthetic data, with some projections suggesting that most available new data for training could be machine-generated by 2030 [1]. This shift to a mainly synthetic content presents a critical challenge: repeated training in synthetic data leads to a phenomenon known as model collapse, where model performance degrades over generations of training, eventually rendering the models ineffective. Although prior studies have explored the causes and detection of model collapse, existing mitigation strategies remain limited.
In this paper, we identify model overconfidence in their self-generated data as a key driver of collapse. Building on this observation, we propose a confidence-aware loss function that downweights high-confidence predictions during training. We introduce a novel loss function we call Truncated Cross Entropy (TCE). We demonstrate that TCE significantly delays model collapse in recursive training.
We provide a model-agnostic framework that links the loss function design to model collapse mitigation and validate our approach both theoretically and empirically, showing that it can extend the model’s fidelity interval before collapse by more than $2.3\times$. Finally, we show that our method generalizes across modalities. These findings suggest that the design of loss functions provides a simple yet powerful tool for preserving the quality of generative models in the era of increasing synthetic data.
@@ 1 Introduction
Generative models have become the foundation for modern AI applications in several modalities, including text, image, code, and audio. Large Language Models (LLMs) such as ChatGPT [2], LLaMA [3], Phi [4], and DeepSeek [5], as well as image generators DALL-E [6] and Imagen [7], all rely on large datasets scraped from the Web. As these models are continuously updated to reflect recent knowledge and linguistic patterns, the need for ever larger and frequently refreshed training corpora has grown substantially. However, this demand is colliding with a shift in the data landscape: synthetic content is increasingly populating the Internet, contaminating the very datasets used for model training.
@@
This shift raises fundamental concerns. Generative models trained on outputs from earlier generations can degrade over time, a phenomenon known as model collapse [8, 9, 10]. Even moderate levels of synthetic data (as little as 1%) have been shown to cause collapse, and scaling the model or the size of the dataset fails to reliably prevent it [11]. Importantly, this is not a hypothetical scenario: recent analysis of the LAION-5B dataset [12], a cornerstone for many generative models, revealed measurable contamination from synthetic sources. [^1] Following prior works [8, 9], we use the term contamination or poisoning to describe the recursive use of synthetic data in training, which leads to degradation in performance over generations. As synthetic data becomes dominant, safeguarding generative model stability becomes an urgent challenge.
@@
Several approaches have been explored to mitigate model collapse. Some studies combine synthetic and real data to slow degradation [13]; others rely on post hoc supervision, such as SIMS [14], using labeled synthetic data as negative signal to improve generations or explore decoding strategies [15]. In practice, however, distinguishing real from synthetic data remains a major challenge. Watermarking methods [16, 17] have been proposed to help identify model-generated content, but they are not widely adopted. More fundamentally, recent critiques [18] highlight that many prior studies simulate collapse in unrealistic settings, such as using purely synthetic training data, limiting the applicability of their conclusions. Despite growing interest in this issue, the role of the loss function, central to how models learn, has received little attention in the context of model collapse.
@@
In this work, we present ForTIFAI, showing that model collapse can be mitigated through principled modifications to training loss. We argue that generative models tend to be overconfident in their synthetic predictions (as depicted in Figure 1), causing feedback loops that degrade performance. This issue is further amplified by approximation errors and finite sampling effects during training, which cause models to concentrate probability mass on a shrinking subset of the true data distribution over generations [9, 10]. To counter this, we introduce Truncated Cross Entropy (TCE), a novel confidence-aware loss function that reduces the weight of high-confidence predictions during training by masking predictions above a confidence threshold. TCE is model-agnostic and easy to implement on top of Cross Entropy (CE), making it attractive for real-world deployment.
@@
  • We introduce a loss-based, model-agnostic method for mitigating model collapse, explicitly addressing sampling bias identified in prior studies.
  • We propose Truncated Cross Entropy (TCE), a novel confidence-aware loss function that downweights high-confidence tokens during training, and demonstrate that it extends model fidelity by more than $2.3\times$.
  • We design a comprehensive evaluation framework grounded in practical and evolving data regimes, including datasets where new real-world content continues to appear, some content is regenerated multiple times while other parts remain untouched, and model quality is measured by retained knowledge rather than token similarity.
  • We demonstrate the generality of our approach by applying it to diverse generative models, including Gaussian Mixture Models (GMMs) and Variational Autoencoders (VAEs), confirming its effectiveness beyond LLMs.
@@ 2 Method
In this section, we first describe the Model Collapse phenomena. We then provide intuition about the correlation between model collapse and token prediction confidence [^2] Throughout this paper, we define a model’s confidence in tokens as the predicted probability assigned to them. in Large Language Models (LLMs). Building upon this intuition, we derive a class of confidence-aware loss functions to thwart autophagy [12] in generative models.
@@ 2.1 Problem Definition
Model Collapse is the gradual degradation that occurs when a model’s own outputs pollute the training data of its successors. Each new model then learns from these possibly flawed examples, causing a drift from the true data distribution [8].
During the generation phase, statistical approximations, limited functional expressiveness, and approximation errors introduce non-zero probabilities of distributional shifts in the model’s outputs.
(a) Confidences over Wikitext
(b) Confidences over re-generated Wikitext
Figure 1: Models display higher confidence in their own generated data when recursively trained. These charts illustrate the probabilities assigned by the LLaMA-3.2-1B model and a version fine-tuned on a fraction of Wikitext. The synthetic dataset in this case is generated by the fine-tuned model. The fine-tuned model shows higher confidence on the portion of the Wikitext it was trained on as well as the entire dataset generated by it. In contrast, the original model demonstrates similar confidence in both cases. This behavioral analysis was key to finding mitigations discussed in the following sections of ForTIFAI.
2.1.1 Key Observation
Figure 1 illustrates a consistent gap in model confidence between self-generated samples and unseen data. We exploit this difference as a signal to identify and discount synthetic samples during training, allowing the model to reduce their influence and mitigate collapse. This gap is a side effect of the approximation error and the greedy sampling performed for generating text [15].
2.2 Confidence-aware Loss Functions
Building on Sec. 2.1.1, we propose confidence-aware loss functions that reduce the influence of overconfident predictions, typically associated with synthetic tokens. By selectively down-weighting, or ignoring, model-generated samples during training using their statistical properties, we mitigate their harmful influence on the loss and reduce the risk of collapse.
2.2.1 Truncated Cross Entropy (TCE)
Building on the findings from Section 2.1.1, we introduce a novel loss function, Truncated Cross Entropy (TCE). In TCE, confident predictions are explicitly masked to remove their influence on training. By selectively masking overly confident predictions, TCE forces the model to learn from less-certain datapoints, thereby preserving distribution tails compared to CE. We formulate TCE as:
@@ 2.2.1 Truncated Cross Entropy (TCE)
2.3 Mathematical Intuition
Figure 2: Effect of TCE on a one-dimensional Gaussian estimator under a fully synthetic training loop. In each generation, $10,000$ samples are generated with sampling bias $\gamma=0.9$. Standard estimation with Cross Entropy (CE) leads to rapid variance collapse. By contrast, TCE with a well-chosen threshold ($\gamma=0.85$) delays collapse substantially. Conversely, if the threshold is set poorly ($\gamma=0.80$), the variance may diverge or converge to zero prematurely.
@@ 2.3 Mathematical Intuition
In a fully synthetic training loop, where each model $G_{t}$ is trained solely on samples from $G_{t-1}$, Alemohammad et al[12] prove that the model’s estimated covariance $\sigma_{t}$ collapses almost surely:
@@
where $\lambda$ represents sampling bias, contracting the variance at each iteration, leading to rapid loss of sample diversity, as illustrated in Figure 2. This decrease in variance is empirically observed across models such as DDPMs [19] and StyleGAN-2 [20].
@@ 3 Model Collapse Evaluation Framework
To realistically assess the impact of Model Collapse, we adopt an experimental framework inspired by [13]. Initially, web data was predominantly human-generated and often curated for quality. Over time, however, machine-generated content, often indistinguishable from human-written text, has increasingly supplemented online data. Our framework simulates a realistic model collapse process by progressively adding synthetic data to the training set. We evaluate the methods on math, reasoning, and knowledge recall benchmark performance. Moreover, we analyze the number of synthetic training iterations it takes for the model to lose performance below a certain threshold, which we call the time to failure.
@@
This setup also follows the conditions mentioned in [18]:
  1. 1.
    Increasing Pre-training Data: Unlike previous studies with fixed dataset sizes, our setup continuously increases the pre-training data at each generation. This growth mirrors real-world trends among state-of-the-art models; for example, LLaMA 1, 2, and 3 were trained on 1.4, 2, and 15 trillion tokens, respectively. Although dataset growth has eventual limits, these bounds remain extremely high (approximately one quadrillion tokens) [21], far exceeding current dataset sizes.
  2. 2.
  3. 3.
@@ 4 Results
To demonstrate the effectiveness of Truncated Cross Entropy, we evaluate it across three settings: Transformers, Variational Autoencoders, and Gaussian Mixture Models. For Transformers, we focus on language modeling, testing LLaMA 3 [3] and Gemma 3 [22] on benchmarks for mathematical reasoning, logical inference, and factual recall. For image generation, we use Variational Autoencoders, and for mixture modeling, we evaluate standard GMMs. Across all experiments, our loss function mitigates model collapse without degrading overall training performance.
@@ 4.1 Experiment Details
We apply the recursive training setup illustrated in Figure 3 on two datasets: Wikitext-2-raw-v1 [23] and a newly curated dataset called Imagination of Web.
@@ 4.1.1 Datasets
The second dataset, Imagination of Web, is a composite dataset constructed from subsets of Wikitext, HellaSwag [24], and GSM8k [25]. It is designed to test the generalization of our loss functions across multiple tasks, such as commonsense, math, and factual recall.
@@ 4.2.1 Knowledge Retention Test
Validation perplexity is a common metric for evaluating LLMs, but it primarily reflects predictive fluency; how well a model guesses the next token in unseen text. While useful for assessing grammar and style, it does not quantify factual knowledge retention. Moreover, low perplexity does not guarantee meaningful or human-like output, as humans do not always select the most likely next word [26, 27].
@@ 4.2.2 Math, Reasoning, and Coherence
To evaluate the impact of model collapse on reasoning and linguistic capability, we use three standard benchmarks: GSM8K [25], Hellaswag [24], and BLiMP [28].
@@
Together, these benchmarks assess the model’s ability to generalize, reason, and maintain coherence under different forms of linguistic stress. For GSM8K and Hellaswag, we use the Language Model Evaluation Harness [29] to ensure consistent evaluation across models.
@@ 4.6 KL Divergence
Figure 7 illustrates that TCE lead to a significantly lower KL divergence over generations compared to Cross Entropy, indicating better preservation of the original data distribution.
(a) LLaMA-3.2-1B
(b) Gemma-3-1b-pt
@@
4.7 Gaussian Mixture Models (GMM)
Unlike deep learning models discussed earlier, Gaussian Mixture Models (GMMs) are not based on neural networks. Instead, they are trained by maximizing the likelihood of the input data under a mixture of $N$ Gaussian components, parameterized by means and variances. GMMs are central to previous theoretical and empirical studies of model collapse [8], and extending our method to this setting underscores its broad applicability.
Our intuition mirrors that of the LLM experiments: we aim to prevent the model from overfitting to high-probability modes while preserving low-probability, diverse samples. To achieve this, we simply exclude data points that fall within the top $\alpha$-percentile of likelihoods under the current model. In our experiments (Figure 8), we set $\gamma=0.8$. This procedure is outlined in Appendix F.
4.8 Variational Auto Encoders (VAE)
To demonstrate the applicability of confidence-aware loss functions beyond language modeling, we extend our analysis to the image generation domain using Variational Auto Encoders (VAEs), following the setup in [8]. We train a VAE on the MNIST dataset [30] using the standard evidence lower bound (ELBO) objective, modified with our truncation function:
@@ 5 Conclusion
In this work, we introduced confidence-aware loss functions, specifically Truncated Cross Entropy as an efficient and simple way of mitigating model collapse. We analyzed TCE’s effect on model collapse under realistic, recursive training conditions. Our evaluations demonstrate that TCE consistently outperforms the widely used Cross Entropy loss in both clean (authentic) and contaminated (synthetic) training regimes.
Our approach is supported by a mathematical formulation that links overconfident predictions to accelerated distributional drift. Empirically, we observed slower growth in KL divergence between generated and original data distributions and extending time to failure, defined as the point of significant performance drop, when using TCE by more than $1.7\times$ compared to Cross Entropy. Through a comprehensive suite of benchmarks, we show that models trained with TCE better retain factual knowledge and generalize more effectively as synthetic data accumulates over generations.
We extended our approach to other generative modeling domains, including Gaussian Mixture Models (GMMs) and Variational Auto Encoders (VAEs). Our results confirm that filtering or down-weighting overconfident predictions can delay or mitigate collapse even in non-neural or image-based generative systems.

### v3

# ForTIFAI: Fending Off Recursive Training Induced Failure for AI Models

@@ Abstract
The increasing reliance on generative AI models is rapidly increasing the volume of synthetic data, with some projections suggesting that most available new data for training could be machine-generated by 2030 Gartner, Inc. (2022). This shift to a mainly synthetic content presents a critical challenge: repeated training in synthetic data leads to a phenomenon known as model collapse, where model performance degrades over generations of training, eventually rendering the models ineffective. While the causes of model collapse are increasingly understood, effective mitigation strategies remain scarce. We address this challenge by leveraging a key insight: auto-regressive models tend to generate text sequences to which they assign high confidence (i.e., high log-likelihood). Based on this observation, we introduce the Truncated-Cross-Entropy (TCE) loss function. TCE mitigates collapse by selectively ignoring high-confidence tokens during training, effectively filtering out likely machine-generated artifacts from the learning process. Our experiments demonstrate that models trained with TCE not only learn effectively but also exhibit significantly increased resilience, tolerating over 2.3× more synthetic data before the onset of collapse. In addition, we provide an open-source benchmark for collapse dynamics in mixed-data settings. Our results demonstrate that confidence-aware training objectives can substantially delay collapse onset, offering a practical and generalizable tool for model robustness under synthetic-data exposure.
@@ 1 Introduction
Generative models have become the foundation for modern AI applications in several modalities, including text, image, code, and audio. Large Language Models (LLMs) such as ChatGPT (OpenAI et al., 2024), LLaMA (Grattafiori et al., 2024) and Gemma (Team et al., 2025), as well as image generators DALL-E (Ramesh et al., 2021) and Imagen (Saharia et al., 2022), all rely on large datasets scraped from the Web. As these models are continuously updated to reflect recent knowledge and linguistic patterns, the need for ever larger and frequently refreshed training corpora has grown substantially. However, this demand is colliding with a shift in the data landscape: synthetic content is increasingly populating the Internet, contaminating the very datasets used for model training.
@@
This shift raises fundamental concerns. Generative models trained on outputs from earlier generations can degrade over time, a phenomenon known as model collapse (Shumailov et al., 2024a; b; Dohmatob et al., 2024b). Even moderate levels of synthetic data (as little as 1%) have been shown to cause collapse, and scaling the model or the size of the dataset fails to reliably prevent it (Dohmatob et al., 2024a). Importantly, this is not a hypothetical scenario: recent analysis of Alemohammad et al. (2023) on the LAION-5B dataset, a cornerstone for many generative models, revealed measurable contamination from synthetic sources. [^1] Following prior works (Shumailov et al., 2024a; b), we use the term contamination or poisoning to describe the recursive use of synthetic data in training, which leads to degradation in performance over generations. As synthetic data becomes dominant, safeguarding generative model stability becomes an urgent challenge.
@@
Several approaches have been explored to mitigate model collapse. Some studies combine synthetic and real data to slow degradation (Gerstgrasser et al., 2024); others rely on post hoc supervision, such as SIMS (Alemohammad et al., 2024), using labeled synthetic data as negative signal to improve generations or explore decoding strategies (Drayson et al., 2025). In practice, however, distinguishing real from synthetic data remains a major challenge. Watermarking methods (Zhang & Koushanfar, 2024; Kirchenbauer et al., 2024) have been proposed to help identify model-generated content, but they are not widely adopted. Moreover, recent critiques (Schaeffer et al., 2025) highlight that many prior studies simulate collapse in unrealistic settings, such as using purely synthetic training data, limiting the applicability of their conclusions. Despite growing interest in this issue, the role of the loss function, central to how models learn, has received little attention in the context of model collapse.
@@
In this work, we present ForTIFAI, showing that model collapse can be mitigated through principled modifications to training loss. We argue that generative models tend to be overconfident in their synthetic predictions (as depicted in Figure 1), causing feedback loops that degrade performance. This issue is further amplified by approximation errors and finite sampling effects during training, which cause models to concentrate probability mass on a shrinking subset of the true data distribution over generations (Shumailov et al., 2024b; Dohmatob et al., 2024b). To address this challenge, we introduce Truncated Cross-Entropy (TCE), a loss function that formalizes the intuition of ignoring what the model already “knows” too well. TCE operates by setting the loss to zero for any token whose predicted probability exceeds a predefined threshold. This prevents the model from overfitting to its own high-conviction and often low-diversity outputs. As a minimal extension of the standard Cross-Entropy (CE) objective, TCE is model-agnostic, computationally efficient, and easy to implement, making it a practical solution for real-world deployment.
@@
  • We establish that a model’s confidence, measured by sequence probability, serves as a strong signal for identifying its own generated text within mixed datasets.
  • We introduce Truncated Cross Entropy (TCE), a confidence-aware loss function that leverages this signal to mitigate model collapse, enabling models to tolerate more than $2.3\times$ greater proportions of synthetic data without degradation.
  • We propose a comprehensive evaluation framework for model collapse that reflects realistic conditions, including the continual accumulation of new data and the need to measure both retained and newly acquired knowledge.
@@ 2.1 Problem Definition
Model Collapse is the gradual degradation that occurs when a model’s own outputs pollute the training data of its successors. Each new model then learns from these possibly flawed examples, causing a drift from the true data distribution (Shumailov et al., 2024a).
The fundamental cause of collapse is the discrepancy between the true data distribution, $P_{data}$, and the model’s learned approximation of it, $P_{\theta}$. Because any realizable model has finite capacity and is trained on finite data, its outputs will contain statistical artifacts not present in the true distribution. Compounding this issue, the decoding process itself introduces a further distributional shift. Common decoding algorithms (e.g., greedy search, beam search) are designed to favor high-probability sequences rather than sampling purely from the model’s learned distribution. This bias means that even if a model were a perfect representation of the data ($P_{\theta}=P_{data}$), the generated text would over-represent the modes of the distribution.
When these skewed outputs are used to train a successor model, the imperfections are learned and amplified. This iterative process leads to a progressive distributional drift, causing a catastrophic loss of information, particularly in the tails of the distribution, ultimately making the model unusable.
Figure 1: A model exhibits significantly higher confidence in its own generated data compared to unseen data. The histograms show token probabilities from a base LLaMA-3.2-1B model (blue) and a fine-tuned version (orange). When evaluating its own synthetic generations (right column), the fine-tuned model shows a clear shift toward higher confidence compared to its evaluation on unseen data (left column). In contrast, the base model’s confidence distribution remains consistent across both datasets, identifying elevated confidence as a sign of self-generated text.
2.2 Key Observation
Figure 1 illustrates a consistent gap in model confidence between self-generated samples and unseen data. We exploit this difference as a signal to identify and remove the effect of synthetic samples during training. This gap is a side effect of the approximation error and the greedy sampling performed for generating text (Drayson et al., 2025).
2.3 Truncated Cross Entropy (TCE)
Building on the findings from Section 2.2, we introduce a novel loss function, Truncated Cross Entropy (TCE). In TCE, confident predictions are explicitly masked to remove their influence on training. By selectively masking overly confident predictions, TCE forces the model to learn from less-certain datapoints, thereby preserving distribution tails compared to CE. We formulate TCE as:
@@ 2.3 Truncated Cross Entropy (TCE)
2.4 Mathematical Insight
@@
In a fully synthetic training loop, where each model $G_{t}$ is trained solely on samples from $G_{t-1}$, Alemohammad et al. (2023) prove that the model’s estimated covariance $\sigma_{t}$ collapses almost surely:
@@
where $\lambda$ represents sampling bias, contracting the variance at each iteration, leading to rapid loss of sample diversity, as illustrated in Figure 2. This decrease in variance is empirically observed across models such as DDPMs (Ho et al., 2020) and StyleGAN-2 (Karras et al., 2020).
Figure 2: Effect of TCE on a one-dimensional Gaussian estimator under a fully synthetic training loop. In each generation, $10,000$ samples are generated with sampling bias $\gamma=0.9$. Standard estimation with Cross Entropy (CE) leads to rapid variance collapse. By contrast, TCE with a well-chosen threshold ($\gamma=0.85$) delays collapse substantially. Conversely, if the threshold is set too small ($\gamma=0.80$), the variance may diverge.
@@ 3 Model Collapse Evaluation Framework
To realistically assess the impact of Model Collapse, we adopt an experimental framework inspired by Gerstgrasser et al. (2024). Initially, web data was predominantly human-generated and often curated for quality. Over time, however, machine-generated content, often indistinguishable from human-written text, has increasingly supplemented online data. Our framework simulates a realistic model collapse process by progressively adding synthetic data to the training set. We evaluate the methods on math, reasoning, and knowledge recall benchmark performance. Moreover, we analyze the number of synthetic training iterations it takes for the model to lose performance below a certain threshold, which we call the time to failure.
@@
This setup also follows the conditions mentioned in Schaeffer et al. (2025):
  1. 1.
    Increasing Pre-training Data: Unlike previous studies with fixed dataset sizes, our setup continuously increases the pre-training data at each generation. This growth mirrors real-world trends among state-of-the-art models; for example, LLaMA 1, 2, and 3 were trained on 1.4, 2, and 15 trillion tokens, respectively. Although dataset growth has eventual limits, these bounds remain extremely high (approximately one quadrillion tokens) Villalobos et al. (2024), far exceeding current dataset sizes.
  2. 2.
  3. 3.
@@ 4 Results
To demonstrate the effectiveness of Truncated Cross Entropy, we evaluate it across three settings: Transformers, Variational Autoencoders, and Gaussian Mixture Models. For Transformers, we focus on language modeling, testing LLaMA 3 Grattafiori et al. (2024) and Gemma 3 Team et al. (2025) on benchmarks for mathematical reasoning, logical inference, and factual recall. For image generation, we use Variational Autoencoders, and for mixture modeling, we evaluate standard GMMs. Across all experiments, our loss function mitigates model collapse without degrading overall training performance.
@@ 4.1 Experiment Details
We apply the recursive training setup illustrated in Figure 3 on two datasets: Wikitext-2-raw-v1 Merity et al. (2016) and a newly curated dataset called Imagination of Web.
@@ 4.1.1 Datasets
The second dataset, Imagination of Web, is a composite dataset constructed from subsets of Wikitext, HellaSwag Zellers et al. (2019), and GSM8k Cobbe et al. (2021). It is designed to test the generalization of our loss functions across multiple tasks, such as commonsense, math, and factual recall.
@@ 4.2.1 Knowledge Retention Test
Validation perplexity is a common metric for evaluating LLMs, but it primarily reflects predictive fluency; how well a model guesses the next token in unseen text. While useful for assessing grammar and style, it does not quantify factual knowledge retention. Moreover, low perplexity does not guarantee meaningful or human-like output, as humans do not always select the most likely next word.
@@ 4.2.2 Math, Reasoning, and Coherence
To evaluate the impact of model collapse on reasoning and linguistic capability, we use three standard benchmarks: GSM8K Cobbe et al. (2021), Hellaswag Zellers et al. (2019), and BLiMP Warstadt et al. (2023).
@@
Together, these benchmarks assess the model’s ability to generalize, reason, and maintain coherence under different forms of linguistic stress. For GSM8K and Hellaswag, we use the Language Model Evaluation Harness Gao et al. (2024) to ensure consistent evaluation across models.
@@ 4.6 KL Divergence
To quantify distributional drift across generations, we measured the Kullback–Leibler (KL) divergence between the token distributions of synthetic datasets and the original WikiText-2 corpus. Each dataset is tokenized with the model’s tokenizer, and normalized token frequencies are used to compute the KL divergence between the original and synthetic distributions.
Figure 7 shows that models trained with TCE maintain a substantially lower KL divergence than those trained with Cross Entropy, indicating better preservation of the original data distribution.
@@ 5 Conclusion
We have demonstrated that a confidence-aware loss function, Truncated Cross-Entropy (TCE), provides a simple and powerful defense against model collapse. Grounded in a mathematical framework linking overconfidence to distributional drift, TCE extends a model’s operational lifetime by over $1.7\times$ compared to standard cross-entropy. Our benchmarks confirm that this approach not only delays collapse but also improves knowledge retention and generalization in environments with accumulating synthetic data, as validated by slower KL divergence growth.
Crucially, the principle of filtering overconfident outputs is not domain-specific. We verified its efficacy in diverse generative systems, including Gaussian Mixture Models and Variational Autoencoders (Appendices F, G). As the line between human and machine-generated data continues to blur, such robust, low-overhead training methods will be essential for building stable and trustworthy AI.